# Patch-wise Graph Contrastive Learning for Image Translation

### 13 Dec 2023 v1

@@ Introduction
In a patch-level view, the image has diverse local semantics. The relational knowledge between the patches embodies the correlation between each region, and is utilized for various image generation tasks.
@@ Results
CycleGAN (Zhu et al. 2017) 77.2 1.957 76.3 3.532 54.6 3.430 84.9 1.022 174.6 10.051
MUNIT (Huang et al. 2018) 133.8 3.790 91.4 6.401 181.7 12.03 115.4 4.901 207.0 12.853
Distance (Benaim and Wolf 2017) 72.0 1.856 81.8 4.410 98.1 5.789 97.2 2.843 181.9 11.362
GCGAN (Fu et al. 2019) 86.7 2.051 105.2 6.824 79.4 5.153 97.5 2.755 178.4 10.828
@@
NEGCUT (Wang et al. 2021) 39.6 0.477 48.5 1.432 51.0 2.338 82.7 1.352 154.1 7.876
LSeSIM (Zheng, Cham, and Cai 2021) 38.0 0.422 49.7 2.867 52.4 3.205 83.9 1.230 168.6 10.386
HnegSRC (Jung, Kwon, and Ye 2022) 34.4 0.438 46.4 0.662 49.2 2.531 81.8 1.181 158.3 8.434
@@
Following the previous work (Park et al. 2020), we verify our method for the single image translation. The input is a Claude Monet’s painting, and the target domain is the natural landscape images. Detailed experimental settings are provided in the supplementary material. For comparison, we choose previous single-image translation models, STROTSS  (Kolkin, Salavon, and Shakhnarovich 2019) and WCT2  (Yoo et al. 2019). Also, we selected the patch-based contrastive learning methods, which are CUT  (Park et al. 2020), FSeSim  (Zheng, Cham, and Cai 2021) and HnegSRC  (Jung, Kwon, and Ye 2022).
@@ Discussion
The proposed method consists of two parts. First, we construct the graph by the pretrained encoder. Second, we utilize top-$K$ pooling by the pooling vector $p$ to focus on task-relevant nodes which provides the localized graph. To investigate the effectiveness of each part, we first investigate what the vector $p$ learns for the graph pooling procedure. Second, we investigate the adjacency matrix $A$ constructed as in Fig. 10 to verify the patch-wise connection used to construct the graph.
@@
Here, we provide the ablation study on the graph, such as the number of hops, number of graph pooling layer, value for similarity threshold, and downsampling ratio of the pooling. First, we provide the ablation study for the number of hops and the similarity threshold for $A$. For the lower ($n=1$) and larger number of hops ($n=3$), we observe that the results are degraded. Also, for both the lowered and increased thresholds ($t=0.0,0.4,0.6$), the results are also degraded from the best setting. Especially in the increased threshold (i.e. sparse connectivity), the model shows much degraded performance. This suggests that a sufficiently dense graph can capture the semantically meaningful topology.
Second, we trained the model with different settings for pooling layers. Without the pooling layer (# of pool=0), the performance degraded as the network do not leverage the information from the focused view. For more pooling layers, the model also shows degraded performance, as the pooled graph has fewer nodes which leads to fewer negative pairs for the contrastive learning. Additionally, we provide the results with varying downsampling rate. For the downsampling of 1/8, the pooled graph consists of fewer nodes, which leads to similar problem with the excessive pooling layers. This again confirms that a sufficiently dense graph after the pooling can capture the semantically meaningful hiearchy. We provide additional ablation study for the graph construction in the supplementary material.
@@ Acknowledgements
This research was supported by National Research foundation of Korea(NRF) (**RS-2023-00262527**)
Ethical Impacts

### 19 Feb 2024 v2

@@ Introduction
In patch-level view, the image has diverse local semantics. The relational knowledge between the patches embodies the correlation between each region, and is utilized for various image generation tasks.
@@ Results
CycleGAN 77.2 1.957 76.3 3.532 54.6 3.430 84.9 1.022 174.6 10.051
MUNIT 133.8 3.790 91.4 6.401 181.7 12.03 115.4 4.901 207.0 12.853
DistanceGAN 72.0 1.856 81.8 4.410 98.1 5.789 97.2 2.843 181.9 11.362
GCGAN 86.7 2.051 105.2 6.824 79.4 5.153 97.5 2.755 178.4 10.828
@@
NEGCUT 39.6 0.477 48.5 1.432 51.0 2.338 82.7 1.352 154.1 7.876
LSeSIM 38.0 0.422 49.7 2.867 52.4 3.205 83.9 1.230 168.6 10.386
HnegSRC 34.4 0.438 46.4 0.662 49.2 2.531 81.8 1.181 158.3 8.434
@@
Following the previous work (Park et al. 2020), we verify our method for the single image translation. The input is a Claude Monet’s painting, and the target domain is the natural landscape images. Detailed experimental settings are provided in the supplementary material. For comparison, we choose previous methods, STROTSS  (Kolkin, Salavon, and Shakhnarovich 2019) and WCT2  (Yoo et al. 2019). Also, we selected the patch-based contrastive learning methods, which are CUT  (Park et al. 2020), FSeSim  (Zheng, Cham, and Cai 2021) and HnegSRC  (Jung, Kwon, and Ye 2022).
@@ Discussion
The proposed method consists of two parts. First, we construct the graph by the pretrained encoder. Second, we utilize top-$K$ pooling by the pooling vector $p$ to focus on task-relevant nodes which provides the localized graph. To investigate the effectiveness of each part, we first investigate what the vector $p$ learns for the graph pooling procedure. Second, we investigate the adjacency matrix $A$ in Fig. 10 to verify the patch-wise connection used to construct the graph.
@@
We provide the ablation study on the graph in Table 2, regarding on the number of hops ($n$), value for similarity threshold ($t$), number of graph pooling layer and downsampling ratio of the pooling. First, for the ablation study on GNN, we observe that results are degraded for both lower ($n=1$) and larger number of hops ($n=3$). Also, our setting shows the best result compared to the other threshold values ($t=0.0,0.4,0.6$). Especially for the increased $t$ (i.e. sparse connectivity), the model shows much degradation. This suggests that sufficiently dense graph captures meaningful topology.
Second, for the ablation study on pooling layers, the performance is degraded without the pooling layer (# of pool=0), as the network do not leverage the information from the focused view. For more pooling layers, models also show degraded performance, as the pooled graphs have fewer nodes which leads to fewer negative pairs for the contrastive learning. Additionally, we provide the results with varying downsampling rate. For 1/8 downsampling, the pooled graph have fewer nodes, which leads to similar problem with the excessive pooling layers. This again confirms that a sufficiently dense graph after the pooling can capture the semantically meaningful hiearchy. We provide additional ablation study for the graph construction in the supplementary material.
@@ Acknowledgements
This research was supported by National Research foundation of Korea(NRF) (**RS-2023-00262527**). This research was also supported by Field-oriented Technology Development Project for Customs Administration through National Research Foundation of Korea(NRF) funded by the Ministry of Science & ICT and Korea Customs Service(**NRF-2021M3I1A1097938**).